### v2

# CoRaiS: Lightweight Real-Time Scheduler for Multi-Edge Cooperative Computing

Abstract
Multi-edge cooperative computing that combines constrained resources of multiple edges into a powerful resource pool has the potential to deliver great benefits, such as a tremendous computing power, improved response time, more diversified services. However, the mass heterogeneous resources composition and lack of scheduling strategies make the modeling and cooperating of multi-edge computing system particularly complicated. This paper first proposes a system-level state evaluation model to shield the complex hardware configurations and redefine the different service capabilities at heterogeneous edges. Secondly, an integer linear programming model is designed to cater for optimally dispatching the distributed arriving requests. Finally, a learning-based lightweight real-time scheduler, CoRaiS, is proposed. CoRaiS embeds the real-time states of multi-edge system and requests information, and combines the embeddings with a policy network to schedule the requests, so that the response time of all requests can be minimized. Evaluation results verify that CoRaiS can make a high-quality scheduling decision in real time, and can be generalized to other multi-edge computing system, regardless of system scales. Characteristic validation also demonstrates that CoRaiS successfully learns to balance loads, perceive real-time state and recognize heterogeneity while scheduling.
Index Terms:
Edge computing, Multi-edge cooperative computing, Deep learning, Real-time scheduling
I Introduction
Edge computing brings computation and storage resources closer to the sources of data, facilitating the processing of client data at the network periphery while meeting stringent response time requirements. Some great progresses have been achieved, especially in the mobile edge computing [1, 2]. However, practical applications often reveal challenges. As shown in Fig. 1, each edge hosts a diverse set of services, and it typically serves multiple clients. The distribution of clients among edges exhibits non-uniformity, with variations in both the number of requests submitted by each client and the specific service they require. This complexity can potentially degrade service quality, especially when an edge is inundated with an excessive number of client requests.
Figure 1: The illustration of unbalanced workloads and resource utilization of edge computing.
The multi-edge cooperative computing combining constrained resources of multiple edges into a powerful resource pool can provide more diversified services and ensure sufficient computing and storage resources. Therefore, it has higher probability to meet the requirements by computation-intensive and latency-critical requests, and improve the average utilization of edges and response time of requests. The edge-cloud system can be regarded as a special case of multi-edge cooperative computing, because the cloud can be considered as an edge with significantly enhanced computational capabilities.
Many researchers are interested in the multi-edge cooperation system, and have proposed some algorithms [3, 4, 5, 6, 7, 8, 9] to schedule independent requests and requests with logical execution order in the system. They usually make some hypothesises to support their research. Firstly, only CPU is configured on the edges, and the number of calculations required to respond to each request is known. With such two conditions, the response time of each request can be obtained by dividing the calculation number by the CPU frequency. Secondly, the request arrival patterns follow known probability distributions, such as poisson distribution, multinomial distribution, etc. With such assumption, many statistical theories and models can be applied to the multi-edge scheduling problems.
However, the assumptions contradict many practical situations. (i) The heterogeneous edges are configured with various computing units, such as CPUs and GPUs. Especially, with increasing requests related to deep learning accessing to edges, it has become a major trend to deploy GPUs at edges. (ii) It is difficult to estimate the calculation number required by requests. Because, when the corresponding processing code is black-box, the calculation number cannot be predicted in advance; when the code is white-box but includes some loop operation and judgment statements, the timing of jumping out of loops and the judgment results are also unpredictable, which makes the calculation numbers impossible to be estimated. (iii) The arrival pattern of requests is almost unpredictable. Each client has its own unique request generation pattern. Though the pattern of a single client can be estimated, it is difficult to analyze the composite pattern of multiple clients. Besides that, another two phenomenons should be focused on, i.e. during the execution process of a service, it will occupy multiple resources such as CPU, GPU, and Memory simultaneously, and the quality of service (QoS) varies across edges [^1] The Quality of Service (QoS) for a specific service exhibits variations among heterogeneous edges due to differences in their hardware configurations. Additionally, even among homogeneous edges with identical hardware configurations, the QoS may differ because they allocate different resources to the service. For instance, if two edges share the same hardware setup, one may assign a 2-core CPU to the service while the other allocates a 4-core CPU. As a result, the QoS of the service will vary between these two edges..
Based on the analysis above, we can identify that the multi-edge cooperative scheduling faces three challenges. (i) Multi-edge cooperative computing system state modeling: The diverse hardware composition and different resource allocation schemes for services at edges pose a severe problem for modeling system state. The QoS of edges would not be fairly evaluated, unless the edges heterogeneity can be shielded and an unified QoS evaluation method can be built. Meanwhile, the system state that includes remaining resource of CPU/GPU/Memory at edges keeps changing as the arrival, running, and completion of requests. Dispatching requests to edges based on the perfect initial state, like traditional approaches, will make the scheduling solution deviate from the optimal at the current state. Therefore, it is challenging but necessary to build a system-level state model that supports edges unified modeling and perceives dynamic changes of resource states. (ii) Multi-edge cooperative scheduling formulation: Most previous theoretical models were designed based on probability distribution assumption of requests and single computing hardware assumption of edges [10, 11], which may not accurately reflect real-world application scenarios. Some learning-based approaches [5, 6] are also studied to optimize the scheduling behaviours on some specific datasets and specially constructed multi-edge network. However, the generalization ability [^2] Here the generalization ability refers to the pre-trained scheduling models can help make effective decisions on unseen applications and multi-edge networks. of the approaches is not good. Once the application and network environment change, a large amount of data (with expert knowledge sometimes) must be collected to retrain the scheduling model, while sometimes it is difficult to collect efficient data and expert knowledge. Therefore, a new mathematical formulation is required to reveal the essence of multi-edge cooperative scheduling problem and to guide scheduling algorithm research for requests with any arrival pattern. (iii) Real-time scheduler designing: The multi-edge scheduling problem for multiple requests is essentially a combinatorial optimization problem with some constraints. The search space of such problem is huge and will grow as the numbers of edges and requests increase. Computing an optimal solution in the huge search space is theoretically time-consuming. The previous works usually took a long time to make the scheduling decision. However, in practice, only methods that support real-time scheduling can ensure the efficient operation of the multi-edge cooperative computing system.
To cope with the inherent challenges mentioned above, we first propose a system-level state evaluation model to express the service-oriented performance and workload of edges at any scheduling period. In this way, the important and differentiated performance characteristics closely related to edge scheduling are preserved, and the heterogeneous configuration of edges can be ignored when making scheduling decisions. Secondly, based on the system-level state evaluation model, we provide a new integer linear programming formulation for the multi-edge cooperation scheduling, which can be a good starting point to inspire solver searching or scheduling algorithms design. Finally, we propose CoRaiS, a reinforcement learning based lightweight real-time scheduler for multi-edge cooperative computing system. Given a high-level goal to minimize the response time of all requests, CoRaiS automatically learns a sophisticated system-level real-time scheduling policy. The policy can be directly generalized to other applications and multi-edge networks.
The main contributions of this paper can be summarized as follows.
  • A system-level state evaluation model is built to capture important features closely related to scheduling across edges, including service-oriented performance feature and workload feature.
  • The multi-edge scheduling problem is presented as a new integer linear programming formulation, which will support the designing and optimization of scheduling algorithms.
  • A lightweight attention-based scheduler (CoRaiS) is proposed to minimize the response time over all requests distributed at edges. The scheduler can provide a high-quality near-optimal solution in real time, irrespective of request arrival patterns and system scales.
The following paper is structured as follows: Section II presents an overview of related work. In Section III, we provide two use cases to illustrate the key aspects of request scheduling in a multi-edge cooperative computing system. Sections IV and V delve into the designed model of the multi-edge cooperative computing system and the formulation of multi-edge scheduling, respectively. Following this, Section VI elaborates on the architecture of the proposed CoRaiS system. Simulation experiments and prototype-based experiments are detailed in Sections VII and VIII, respectively. Finally, the conclusion in Section IX summarizes the findings and contributions of the paper.
II Related Work
Previous studies have delved into the formulation and optimization of scheduling problems within multi-edge or edge-cloud cooperative computing systems, yielding significant contributions [12, 13, 14]. Typically, these works categorize dispatched requests into two groups: independent requests and requests with a logical execution order, subsequently designing specific scheduling algorithms for each category.
II-A Independent Requests Scheduling
Numerous efforts [15, 4, 5, 6, 16, 10, 11, 17, 18] have primarily focused on independent requests. Most of them formulate the request scheduling problem within multi-edge cooperative system as integer linear programming problem, and make some assumptions to enable the formulations. For instance, [15, 4] assume known computation requirements and average arrival rates, formulating the problem as an integer linear program and proving its NP-hardness. They further propose heuristic approximation algorithms based on linear program relaxation and rounding to address the issue. Other works like [10, 11] make assumptions about request arrival patterns and computation requirements to address scheduling challenges. The works [17, 18] assume weightings indicating latency sensitivity and machine-dependent processing times for requests and propose online scheduling frameworks to minimize total weighted response time across all requests. The work [16] formulates a request routing problem under the assumption of computation capacity in multi-cell mobile edge computing networks, proposing a randomized rounding algorithm to minimize load to the cloud.
Besides the heuristic methods, learning-based approaches have emerged as a significant contributor to addressing multi-edge cooperative computing challenges. For instance, the work [5] develops its dedicated experimental environment comprising physical servers and leased clouds, introducing a coordinated multi-agent actor-critic algorithm for decentralized request dispatch. This algorithm is evaluated using real-world workload traces from Alibaba. Similarly, EdgeMatrix [6] presents a learning-based scheduling model aiming to maximize throughput while ensuring various Service-Level-Agreement priorities. A networked multi-agent actor-critic algorithm is proposed to customize resource channels and enhance system stability. The work [19] proposes a task scheduling framework across edge clouds, namely LsiA3CS, which employs A3C [20] and heuristic guidance to achieve distributed, asynchronous task scheduling for large-scale IIoT. Aiming at optimizing the average response time of task scheduling and the average energy consumption of the system, a multi-objective task scheduling model is designed, and a task scheduling policy optimization algorithm based on improved asynchronous advantage actor-critic (A3C) is proposed in [21].
Additionally, the optimization objective of researches are also varying. The work [22] proposes a fine-grained two-sided market via an intercloud broker, named SkyPilot to allow users to view the cloud ecosystem and schedule requests in an economic approach. The work [23] uses mixed integer programming to formulate the scheduling problem, which aims to satisfy the maximum number of requests given their deadline requirements. Moreover, the work presents a heuristic approach using the genetic algorithm to generate the scheduling decision. In the work [24], an optimal control theory (OCTS) based method is introduced. It aims to achieve global optimal results, ensuring computational and transmission latency requirements are met while load balancing various vehicle tasks. The work [25] proposes an evolutionary genetic (GA)-based optimization algorithm, aiming to strike a balance between task execution time and processing cost. The work [26] presents a Dynamic Parallel Computing Offload and Energy Management (DPCOEM) algorithm. It utilizes Lyapunov optimization technology to minimize task response time and achieve near-optimal performance. The work [27] formulates a bilevel optimization problem and proposes ant colony system and monotonic optimization method to minimize the energy consumption of all requests with deadline. The work [28] introduces a resource-constrained task scheduling profit optimization algorithm (RCTSPO). It focuses on profit maximization while enhancing task scheduling time, reliability, and load balancing.
TABLE I: Synthesis table of related works
Related works Computing resources Assumptions of request arrival pattern Assumptions of execution time for each request Problem formulations Optimization objectives Scheduling algorithms
[15, 4] CPU The average arrival rate at time slots is known N/A Mixed integer linear program (MILP) Maximize the number of requests served per slot Greedy algorithm and linear program
[10] CPU The arrival pattern at each edge node follows a Poisson process Requests’ CPU cycles follow exponential distribution. The computing time is estimated via FDC, i.e. the maximum frequency of edges divides required CPU cycles. Mixed integer non-linear programming problem (MINonILP) Minimize the service response time and outsourcing traffic to central clouds Iterative optimization algorithm
[11] CPU The average arrival rate at time slots is known Requests’ CPU cycles are randomly sampled with a finite expection. The computing time is estimated via FDC. Queuing game Minimize the expected cost of each individual edge server Cooperative queueing game approach
[17, 18] CPU N/A Requests’ CPU cycles are known and an edge can execute at most one request at a time Dual integer programming Minimize the total weighted response time of all requests Online algorithm with dispatching policy and the scheduling policy
[23] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. MILP Maximize the number of requests with deadlines Genetic algorithm
[26] CPU The arriving probability at each slot is known Requests’ CPU cycles are known, and the computing time is estimated via FDC. MINonILP Minimize response time and packet losses of tasks under the limitation of energy queue stability Lyapunov-based dynamic parallel computing offloading
[25] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. MILP Achieve a trade-off between execution time and monetary cost to complete requests Particle swarm optimization
[27] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. MILP Minimize the total energy consumption of all mobile users under the delay constraint Ant colony and monotonic optimization
[28] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. Petri Net Load balancing and profit optimization Heuristic algorithm
[5] CPU Specified based on dataset Obtained from dataset Markov game formulation Balance the workloads among edge and offload some requests to the cloud Coordinated multi-agent actor-critic
[6] CPU Specified based on dataset Obtained from dataset Markov game formulation and MILP Maximize the throughput while guaranteeing various SLA priorities Multi-agent actor-critic algorithm
[19] CPU Specified based on dataset Obtained from dataset Markov game formulation Maximize the long-term throughput rate Actor-critic algorithm
[29] CPU Stochastic requests arrivals The average execution time is known (from dataset) Markov decision process (MDP) Minimize average request completion time Policy gradient
[30] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. Search problem Maximize system throughput Graph-aware encoder-edecoder reinforcement learning model
[31] CPU N/A Requests’ CPU cycles are known, and the computing time is estimated via FDC. MDP Minimize the long-term average delay and energy consumption, and maximize the system throughput Value decomposition multi-agent deep Q learning
[32, 7, 8] GPU and CPU N/A Measured in the hardware environment MDP Minimize the training time of a deep neural network Reinforcement learning (RL)
[33] GPU and CPU N/A Measured in the hardware environment MILP and MDP Minimize the training time of deep neural networks Multi-agent RL
Ours Service-oriented any resource combinations N/A Estimated through the proposed system-level state evaluation model MILP Minimize the average response time of computing requests RL
II-B Scheduling Requests with Logical Execution Order
Dispatching requests with logical execution orders constitutes an important area of study within multi-edge cooperative computing. The dependencies among such requests are often represented as directed acyclic graphs (DAGs), leading to what is commonly referred to as the DAG scheduling problem.
The work [29] introduces Decima, which devises novel representations for requests’ dependency graphs. It employs scalable reinforcement learning models to learn workload-specific scheduling algorithms autonomously, aiming to minimize average request completion time. In [30], computing devices are allocated to continuous data flows within a large distributed system. The paper presents a graph-aware encode-decoder framework designed to learn a generalizable resource allocation strategy. The work [31] proposes an online concurrent user request scheduling mechanism based on multi-agent deep reinforcement learning to optimize the long-term average delay and energy consumption.
In addition to traditional request scheduling, the optimization of device placement for decomposed deep neural networks (DNNs) has also garnered attentions in the multi-edge cooperative scheduling domain, inspired by advancements in neural network-based intelligent large models. Studies also treat the architecture of a DNN model as a DAG, with device placement specifying how each operation in the model should be matched to networked heterogeneous CPU and GPU devices. For example, Mirhoseini et al. [32] propose a sequence-to-sequence model that optimizes device placement for TensorFlow computational graphs. Spotlight [7] is devised to determine optimal device placement for DNN training. It formulates the problem as a Markov decision process with multiple stages and employs a novel reinforcement learning algorithm based on proximal policy optimization. A two-level hierarchical model [8] is also introduced for device placement in neural networks containing tens of thousands of operations. The work [33] proposes TapFinger, a distributed scheduler for edge clusters that minimizes the total completion time of ML tasks through co-optimizing task placement and fine-grained multi-resource allocation.
II-C Related Work Analysis
We analyze the related works from perspectives of computing resources, assumptions on request arrival patterns and computation requirements, problem formulations, optimization objectives and scheduling algorithms, and summarized them in Table I. Delving into the details of previous researches, we can find that previous research often made assumptions based on simplified scenarios, such as only having CPUs on the edges and knowing the number of calculations required for each request. However, in reality, heterogeneous edges equipped with various computing units, including GPUs and TPUs, are becoming increasingly common. Moreover, with the rise in requests related to deep learning, deploying GPUs at the edges has emerged as a significant trend. This shift introduces a new layer of complexity. The interaction of multiple computing resources (CPU, GPU, TPU, etc.), each with its own capabilities and characteristics, coupled with the intricate execution logic of intelligent applications, poses challenges in accurately estimating the running time of requests. Additionally, many works in the field make assumptions about request arrival patterns following known probability distributions like Poisson or multinomial distributions. However, the reality is that the arrival pattern of requests is often unpredictable. Each client may indeed have its own unique request generation pattern, making it challenging to analyze the composite pattern of multiple clients.
In contrast to existing works, our research liberates from many assumptions. Firstly, we extend computing devices to encompass a combination of CPUs and GPUs, facilitating the development of complex intelligent applications. We achieve this by constructing a novel system-level state evaluation model capable of evaluating the computation time of each request. This model is compatible with heterogeneous hardware configurations and diverse intelligent applications. Secondly, we do not constrain the request arrival pattern. Instead, we design a new multi-edge cooperative framework and propose a corresponding lightweight scheduling algorithm to enable real-time scheduling of requests. Both the framework and algorithm are designed to accommodate any arrival pattern of requests.
III Motivating Use Cases
Figure 2: An illustration of the motivating use case.
Figure 3: An illustration of the motivating use case with mobility.
We present two use cases to illustrate the key points of request scheduling in multi-edge cooperative computing system. As shown in Fig. 2, five edges coordinate to serve clients, but the service capability and system status of edges are different, i.e., the hardware and software configure may be heterogeneous, and the backlogs and deployed service replicas can be different as well. Meanwhile, clients have different number of requests that need to be processed. Fig. 3 considers client mobility into the use case, i.e. clients can move to another client region after submitting their requests. Request scheduling has to dispatch each request to an edge and make the total response time of requests minimized with consideration of the edge heterogeneity and client mobility.
Multi-edge cooperative computing can also offer significant benefits in the context of content delivery networks (CDNs) [34, 35, 36], which are distributed networks of servers strategically positioned worldwide to deliver web pages, images, videos, and digital assets with high performance and availability. For instance, in video streaming applications, traditional CDN setups involve caching video content at edge servers for faster delivery. However, video content often requires transcoding into different formats or bitrates to suit various devices and network conditions. Multi-edge computing enables real-time video transcoding at edge servers. When a user requests a video, the edge server dynamically transcodes it into the appropriate format or bitrate based on the user’s device and network conditions. This eliminates the demands for centralized transcoding servers and reduces latency by delivering optimized video directly from the edge server. Similarly, in online gaming, CDNs optimize content delivery by caching game assets, updates, and patches at edge servers. Multi-edge computing can enhance this by distributing these assets across multiple edge locations, ensuring faster downloads and updates for players. Moreover, edge servers can dynamically adjust content delivery based on players’ location, network conditions, and device capabilities, ensuring optimal performance for each player.
However, there are two challenges when realizing the real-time effective scheduling. Firstly, there is a lack of effective business models or interoperability models to facilitate the cooperative interaction among edges and the precise execution of the scheduling decision. Because the traditional cloud business pattern, which submits all requests to the cloud management platform and lets the platform schedule the requests to specific servers, suffers from the huge data traffic consumption, the platform bottleneck and longer response time. Collaboration based on distributed consensus can lead to long-term negotiations at the edges, resulting in delayed scheduling decisions. Secondly, the scheduling space is $N^{M}$, where $N$ and $M$ refer to the number of edges and requests, respectively, indicating that the solution space will increase exponentially with the growth of edge numbers and request numbers. The huge scheduling space makes it difficult to search the optimized solution in time, while the practical scenario requires immediate solutions to enable the multi-edge system operation.
In the following, we will describe our interoperability models for the multi-edge cooperative computing system and formulate the scheduling problems in detail.
IV Multi-edge Cooperative Computing System Model
IV-A Service-oriented Multi-edge Computing System Modeling
Figure 4: Multi-edge cooperative computing system modeling. Part A: The edges are physically interconnected. Each edge has its dedicated client region and the services deployed on each edge operate independently. Part B: The multi-edge cooperative computing system is decomposed into multiple independent service-oriented subsystems, where the edges with same service are grouped.
The multi-edge cooperative computing system consists of a set of network edges =$\mathcal{E}=\{e_{n}\}_{n=1}^{N}$, and can provide diverse services $\mathcal{S}=\{s_{k}\}_{k=1}^{K}$, as shown in Fig. 4. To enable service-oriented effective scheduling, we decompose the multi-edge cooperative system into multiple service-oriented subsystems, i.e. $\{SR_{k}\}_{k=1}^{K}$, where $SR_{k}=\{e_{k1},e_{k2},...,e_{kn^{\prime}}\}$, $e_{ki}$ denotes the $i^{th}$ edge with service $s_{k}$. With the decomposition, our focus can then shift towards addressing scheduling problem within each individual subsystem, since the process of scheduling over the whole multi-edge system can be decomposed into scheduling on each $SR_{k}$. For the sake of clarity, in the following discussions, we omit explicitly declaring the $s_{k}$-oriented. However, it’s important to understand that all the described operations can be applied into any service-oriented subsystems.
Figure 5: Territorial schedulers are introduced into the service-oriented multi-edge cooperative computing system to facilitate effective and real-time scheduling.
We further introduce territorial schedulers into the service-oriented multi-edge cooperative computing system. The novel multi-edge cooperative architecture is presented in Fig. 5. All edges are networked, which enabling data transmission between any two edges. Each edge is under the exclusive management of a single scheduler, which decides the execution location of each received request in its control region. Moreover, every edge within the system willingly collaborates with others and entrusts its scheduler’s decision-making efficacy. Consequently, each edge actively processes all requests dispatched by the scheduler. This collective commitment ensures a seamless and efficient process, contributing to the overarching goal of achieving timely completion of all tasks.
In addition, when an edge device seeks to join the system and contribute its computational resources, it must actively establish a connection with a base station. Subsequently, it is required to transmit registration messages to the territorial scheduler. These registration messages should include details such as deployed services, the number of corresponding replicas, computational capabilities, and geographical locations. Following the registration process, the edge device is expected to periodically send its status information to its designated scheduler, indicating its availability to handle incoming requests. If the information fails to be received continuously for a predetermined threshold number of times, the scheduler will infer that the edge device has gone offline, then if the edge wants to join the system again, it has to register like a new edge. By performing these operations, clients have the capability to submit requests to the edge device through either wireless or wired communication channels. Simultaneously, the scheduler is empowered to efficiently dispatch requests to the appropriate edge devices
IV-B Interoperability Model of Multi-edge Cooperative Scheduling
IV-B1 Concept definition
Some concepts used in the interoperability model of multi-edge cooperative scheduling are described as follows.
Request: A request typically comprises description text and physical input data. The description text outlines essential details such as the required service, client ID, source edge location from which the request originated, and the predicted edge location where the client will be located upon request completion. Input data refers to the actual data intended for processing. For instance, in an image classification request, the description text would specify the required classifier, while the related images would be uploaded as input data
Request brief: A request brief is essentially a data package containing only descriptive information about the request, devoid of detailed input data. For instance, in the context of an image classification request, the brief would encompass details such as the image sizes and the required classifier, while omitting the actual image content. Additionally, the brief includes information about the source edge location and the predicted edge location. The introduction of request briefs leads to a significant reduction in the size of data packages used for scheduling decisions. This reduction in package size has a direct impact on communication delay among edges and the scheduler, resulting in improved efficiency.
Scheduling decision: It contains the information which edge will execute the requests. The edges are required to adhere to this decision, either responding to the requests locally or transferring them to other edges as directed.
IV-B2 Interoperability scheduling model
Based on the architecture shown in Fig. 5, we further propose an interoperability scheduling model to enable the effective multi-edge cooperation.
The multi-edge cooperative scheduling processes involve eight steps: (i) Clients submit requests and related data to edges, along with their predicted edge location after request completion. (ii) Edges receive requests and generate a request brief for each. (iii) Edges provide their schedulers with current service capacity information and the request briefs. (iv) The scheduler makes scheduling decisions based on edge status, request briefs, and designated scheduling algorithms. (v) The scheduler informs edges of the decision, specifying the execution edge for each request. (vi) Based on the decision, edges locally handle the requests or transfer them to other edges along with the relevant data. (vii) The execution edges provide computing results feedback to the predicted edges. (viii) The predicted edges transmit the computing results to clients.
Figure 6: An example of multi-edge cooperative scheduling considering client mobility within a single control region. In Scheduler A’s control region, the submitted request is executed locally, indicating the source edge and the execution edge are the same. In Scheduler B’s control region, the request is transferred to another edge for processing. In this scenario, Scheduler B must simultaneously inform both the source edge and the execution edge (actions 5.1 and 5.2). Subsequently, the source edge transmits all relevant data of the request to the execution edge (action 6.1), which then follows the scheduling decision and handles the request accordingly (action 6.2). Once the request is completed, the execution edge transmits the results to the prediction edge (action 7), which in turn provides feedback to the client (action 8).
Figure 7: In a scenario involving multi-edge cooperative scheduling and client mobility across multiple control regions, schedulers are limited to scheduling edges within their respective control regions. When a request extends beyond this boundary, the scheduler dispatches it to one of its managed edges. Upon completion of the request, the execution edge transmits the results to the predicted edge via cross-region communications (action 7).
To provide a clearer illustration of the interaction processes between schedulers and edges supporting effective request response, we present two examples shown in Fig. 6 and Fig. 7. These examples encompass scenarios involving client mobility both within the control region of a scheduler and beyond its boundaries.
V Multi-edge Scheduling Formulation
In this section, we will formulate the multi-edge scheduling problem in detail.
V-A System-level State Evaluation Model
Figure 8: The linear relationship between running time and the size of data packets for Model Soups [37] on heterogeneous devices. Points are actual statistical data, while lines are function relationships fitted based on the data.
The system-level state evaluation model is inspired by some observations. The first key observation is that there are many popular artificial intelligence services, such as image classification and object detection, exhibiting a functional relationship between their response time and the size of the data packets to be processed. We select the most advance image classification model Model Soups [37] as an example to illustrate the observation, and visualize the relationship of Model Soups in Fig. 8. In Fig. 8, we can observe that the running time and data size of Model Soups is linearly correlated, and the coefficients of the linear function are related to the configuration of computing devices. The most popular object detection model, YOLOv6 [38, 39] also cares about the computation efficiency. They reports the linear relationship between running time and data size for different vision of YOLOv6 when inferring on Tesla T4. Please note, in this paper, we refer to services that have a functional relationship between runtime and data size as ideal services, and our study primarily focuses on this category. The second observation is that many advanced technologies, such as Docker and Kubernetes, empower one service to create multiple independent replicas on a single device and enable the service to specify the required resources. This reservation mechanism ensures that the resources allocated to each service replica are safeguarded against preemption by other processes, thereby stabilizing the QoS of the service.
Based on these observations, we build our system-level state evaluation model. The model consists of two parts: service-oriented performance estimation and service-oriented workload evaluation.
V-A1 Service-oriented Performance Estimation
Two indicators are used to consistently evaluate the service-oriented performance of edges, that are computation time estimation function ($\phi(x)$) and service replica numbers ($\zeta$). $\phi_{i}(x)$ is a function that depicts the relationship between the response time and the size of data packets to be processed at edge $e_{i}$, with $x$ denotes the size of the data packets. $\phi_{i}(x)$ can be approximated by fitting the relationship between data volume and the actual processing time of historical requests, like the fitting operation in Fig. 8. There are some tools that can help establish the relationship, such as numpy.polyfit and scipy.optimize.curve_fit. The hardware configuration also has a significant impact on the execution efficiency, causing $\phi(x)$ to vary across different edges. Fig. 8 illustrates the impact as well. Therefore, when establishing the relationship, only local historical data can be selected. $\zeta_{i}$ refers to the replica number of the service on edge $e_{i}$, which is a predefined system-level parameter to support service parallels. A larger $\zeta_{i}$ indicates edge $e_{i}$ can deal with more requests in parallel.
By defining $\phi(x)$ and $\zeta$, we can focus on considering the primary performance factors of edges while formulating multi-edge cooperative scheduling problem and designing algorithms, and ignore the secondary factors (the heterogeneous hardware configuration and various resource allocation mechanism to services at edges). Furthermore, with $\phi_{i}(x)$, we can predict response times on any edges by inputting the request’s data size. This eliminates the need to analyze black box or white box code to obtain the necessary computation numbers for the request and restricts the computing device to the CPU.
V-A2 Service-oriented Workload Evaluation
Figure 9: Illustration of the state transition of requests across queues. Clients submit requests to $Q^{r}$. After receiving the scheduling decision, requests that will be locally executed are transferred from $Q^{r}$ to $Q^{le}$, and requests that will be executed on other edges are transferred from $Q^{r}$ to $Q^{out}$. The requests that are transferred from other edges are saved in $Q^{in}$. Once the related data of requests in $Q^{in}$ are received, the requests are transferred from $Q^{in}$ to $Q^{le}$. If the requests are completed, then they are transferred from $Q^{le}$ to $Q^{F}$.
We design service-oriented queue models for edge $e_{i}$ to save requests that are at different status, including $Q_{i}^{r}$ for requests that are waiting for scheduling, $Q_{i}^{le}$ for requests that will be executed locally , $Q_{i}^{out}$ for requests that will be transmitted to other edges, $Q_{i}^{in}$ for requests that will be transferred in from other edges and $Q_{i}^{F}$ for requests that have been completed. The state transition of requests across queues is presented in Fig. 9.
When evaluating the workload of edge $e_{i}$, we focus on the requests in $Q_{i}^{le}$ and $Q_{i}^{in}$, because only requests in these two queues will make use of the local resources of edge $e_{i}$. Five features are introduced to evaluate the workloads, including required computing time to complete requests in $Q_{i}^{le}$ (referred to as $c_{i}^{le}$), required data transmission time for requests in $Q_{i}^{in}$ (referred to as $t_{i}^{in}$), required computing time to complete requests in $Q_{i}^{in}$ (referred to as $c_{i}^{in}$), required results transmission time for requests in $Q_{i}^{le}$ (referred to as $b_{i}^{le}$) and required results transmission time for requests in $Q_{i}^{in}$ (referred to as $b_{i}^{in}$).
The five features can be evaluated in customized mathematical approximation models. In this paper, we compute them by (1)-(5) respectively. $r$ refers to a request. $f_{r}$ refers to the data size of request $r$ and $u_{r}$ denotes data volume of $r$’s results. $\varpi_{r}$, $e_{q}$ and $\delta_{r}$ represent the source edge, the execution edge and the predicted edge of request $r$, respectively. $\psi(e_{q},\varpi_{r})$ computes the distance between $e_{q}$ and $\varpi_{r}$, while $\psi(e_{q},\delta_{r})$ obtains the distance between $e_{q}$ and $\delta_{r}$. $C_{t}$ is a constant to represent the transmission speed for unit data through unit distance.
$c_{i}^{le}=\frac{\sum_{r\in Q_{i}^{le}}\phi_{i}(f_{r})}{\zeta_{i}}$ (1)
$t_{i}^{in}=\max\limits_{r\in Q_{i}^{in}}C_{t}f_{r}\psi(e_{q},\varpi_{r})$ (2)
$c_{i}^{in}=\frac{\sum_{r\in Q_{i}^{in}}\phi_{i}(f_{r})}{\zeta_{i}}$ (3)
$b_{i}^{le}=\max\limits_{r\in Q_{i}^{in}}C_{t}u_{r}\psi(e_{q},\delta_{r})$ (4)
$b_{i}^{in}=\max\limits_{r\in Q_{i}^{in}}C_{t}u_{r}\psi(e_{q},\delta_{r})$ (5)
In (1) and (3), we average the computation time required to complete all tasks across multiple copies. As for (2), (4) and (5), we make two assumptions based on experience to predict required data transmission time. Firstly, the data transmission time is positively correlated with both data size and transmission distance. Secondly, edges can simultaneously receive data sent by other edges from different ports.
Evaluating the workload before each scheduling operation allows us to obtain the real-time service capacity of edge $e_{q}$. This real-time system state knowledge is instrumental in making well-informed scheduling decisions.
TABLE II: The definitions of primary notations
Notation Definition
$SR_{k}$ The service $s_{k}$ oriented multi-edge cooperative subsystem.
$r_{z}$ The request $z$
$e_{q}$ The edge $q$
$s_{k}$ The service $k$
$l_{zq}$ Whether $r_{z}$ is located at $e_{q}$. $l_{zq}\in\{0,1\}$
$x_{zq}$ Whether $r_{z}$ is dispatched to $e_{q}$. $x_{zq}\in\{0,1\}$
$f_{z}$ The input data size of $r_{z}$.
$u_{z}$ The output data size of $r_{z}$.
$\phi_{q}(f_{z})$ The computation time to deal with data of $r_{z}$ at $e_{q}$.
$p_{q}$ The replica number of $s_{k}$ at $e_{q}$
$c_{q}^{le}$ The computation time to complete backlogs in $Q_{q}^{le}$.
$c_{q}^{in}$ The computation time to complete backlogs in $Q_{q}^{in}$.
$\psi(e_{q},e_{n})$ The transmission distance $e_{n}$ and $e_{q}$.
$C_{t}$ A constant to represent the transmission speed for unit data through unit distance.
$t_{q}^{in}$ The remaining transmission time of backlogs in $Q_{q}^{in}$.
$[Z]$1, 2, …, Z}. Z denotes the number of requests.
$[Q]$1, 2, …, Q}. Q denotes the number of edges in one control region.
$[N]$1, 2, …, N}. N denotes the number of edges in the whole $SR_{k}$.
V-B Problem Formulation
Each service $s_{k}$ oriented subsystem $SR_{k}$ has multiple control regions. Each control region can be formulated as $CoMEC=(\mathcal{E},\mathcal{V},\mathcal{P},\mathcal{I})$. =$\mathcal{E}=\{e_{q}\}_{q=1}^{Q}$ is a set of network edges where $s_{k}$ is deployed, $|\mathcal{E}|=Q$ is the number of edges. $\mathcal{V}=\{\phi_{q}\}_{q=1}^{Q}$ is a set of functions and $\phi_{q}$ represents computation time estimation function of $e_{q}$ (explained in the section V-A1). $\mathcal{P}$ is a $Q$-dimensional vector as well. $p_{i}$ denotes the replica numbers of $s_{k}$ at $e_{i}$. $\mathcal{I}$ is a $N\times 3$ matrix that specifies the current workload evaluation of edges. The entry of $\mathcal{I}$ at $q^{th}$ row, denoted as $I_{q}$, represents the workload evaluation of $e_{q}$. $I_{q}=(c_{q}^{le},c_{q}^{in},t_{q}^{in})$ (the definitions are explained in the section V-A2). Additionally, we define $N$ to represent the number of edges in the whole $SR_{k}$. $\psi(e_{q},e_{n})$ is a function to obtain the data transmission distance between $e_{q}$ and $e_{n}$.
The requests distributed in an individual $CoMEC$ can be modeled as $CoR=(\mathcal{R},\mathcal{L},\mathcal{H},\mathcal{F},\mathcal{D})$. =$\mathcal{R}=\{r_{z}\}_{z=1}^{Z}$ is a set of requests that require $s_{k}$, $|\mathcal{R}|=Z$ is the number of requests. $\mathcal{L}$ is a $Z\times Q$ matrix, the entry $l_{zq}$ represents whether $r_{z}$ is located at $e_{q}$ before scheduling, if yes, $l_{zq}=1$, else, $l_{zq}=0$. $\mathcal{H}$ is a $Z\times N$ matrix, the entry $h_{zn}$ represents whether the client who submits $r_{z}$ will move to the client region of $e_{n}$ after $r_{z}$ is completed. If yes, $h_{zn}=1$, else, $h_{zn}=0$. =$\mathcal{F}=\{f_{z}\}_{z=1}^{Z}$ is a set that records the size of input data for all $r_{z}\in\mathcal{R}$. While $\mathcal{U}=\{u_{z}\}_{z=1}^{Z}$ is a set that records the size of results for all $r_{z}\in\mathcal{R}$. $\mathcal{D}$ keeps the practical related data of requests.
Given $CoMEC$ and $CoR$, let $\mathcal{X}\in\{0,1\}^{Z\times Q}$ be a permutation matrix. For $\forall$ $x_{zq}\in\mathcal{X}$, $x_{zq}=1$ represents $r_{z}$ is dispatched to $e_{q}$. Then the objective function of multi-edge cooperative scheduling problem can be formulated as (6). The parameter $T_{q}$ in (6) denotes the required time to complete all requests that are scheduled to $e_{q}$ as $\mathcal{X}$. (6) indicates that the purpose is to get a $\mathcal{X}$ that can minimize the response time over all edges and all requests. $T_{q}$ in (6) can be computed by (7)-(13) with constraints (14) and (15). The definitions of primary notations in the formulation are summarized in Table II.
$obj.$ $~{}~{}\min\limits_{\mathcal{X}}~{}\max\limits_{e_{q}\in\mathcal{E}}~{}~{}~{}T_{q}$ (6)
$\mu_{q}=\frac{\sum_{z\in[Z]}l_{zq}x_{zq}\phi_{q}(f_{z})}{p_{q}}+c_{q}^{le}$ (7)
$\eta_{q}=\frac{\sum_{z\in[Z]}(1-l_{zq})x_{zq}\phi_{q}(f_{z})}{p_{q}}+c_{q}^{in}$ (8)
$\upsilon_{q}^{1}=\max\limits_{z\in[Z]}~{}x_{zq}f_{z}(\sum_{m\in[Q]}l_{zm}\psi(e_{m},e_{q}))$ (9)
$\kappa_{q}^{1}=\max(C_{t}\upsilon_{q}^{1},t_{q}^{in})$ (10)
$\upsilon_{q}^{2}=\max_{z\in[Z]}x_{zq}u_{z}(\sum_{n\in[N]}h_{zn}\psi(e_{q},e_{n}))$ (11)
$\kappa_{q}^{2}=\max(C_{t}\upsilon_{q}^{2},b_{q}^{in})$ (12)
$T_{q}=\max(\kappa_{q}^{1},\mu_{q})+\max(\eta_{q}+\kappa_{q}^{2},b_{q}^{in})$ (13)
$s.t.~{}$ $\sum_{q\in[Q]}x_{zq}=1,\forall z\in[Z]$ (14)
$x_{zq}\in\{0,1\},\forall z\in[Z],\forall q\in[Q]$ (15)
To be specific, (7) predicts the required computing time to complete all requests that will be executed locally, including previous backlogs and new requests scheduled as $\mathcal{X}$. (8) evaluates the required computing time to complete all requests that are transferred from other edges to $e_{q}$, taking into account both backlogs and new requests scheduled as $\mathcal{X}$. (9) predicts the required longest transmission time for data of new requests that are transferred from other edges to $e_{q}$ as $\mathcal{X}$, since multiple edges transmitting data to the same edge in parallel is allowed in our multi-edge cooperative computing system. Considering that the data of backlogs may haven’t been transmitted to $e_{q}$, we design (10) to estimate the maximum transmission time for all requests that are transferred from other edges to $e_{q}$, including both backlogs and new requests scheduled as $\mathcal{X}$, through $\max$ operation over $C_{t}v_{q}$ and $t_{q}^{in}$. (11) and (12) are used to estimate the longest transmission time from the execution edge $e_{q}$ to the predicted edge $e_{n}$ for requests that are processed at $e_{q}$, including new requests and backlogs. Furthermore, due to data transmission and local computing can run simultaneously, and the requests in $Q_{q}^{in}$ cannot be executed until all relevant data has been received by $e_{q}$, we design (13) to predict the required response time to complete all requests that are scheduled to $e_{q}$ as $\mathcal{X}$, including backlogs and new requests as well. Here $\max$ operation is used to select the case that consumes longer time. While computing a feasible $\mathcal{X}$, constrains (14) and (15) must be met to ensure all requests can be scheduled to only one available edge.
Throughout the system’s operation, there will be multiple task scheduling rounds. In each scheduling round, the formulation can be utilized to obtain the optimal solution $\mathcal{X}$ based on the current system state $CoMEC$ and request state $CoR$.
VI Lightweight Real-time Scheduler Design
The scheduling problem has been represented as an integer linear programming formulation. There have been some solvers, such as Gurobi and Cplex, which can accurately solve such problems. However, the search space of multi-edge cooperative scheduling problem is $Q^{Z}$, and it will significantly grow as the number of edges or requests increases. Therefore, getting the optimal solution is theoretically time-consuming. It is necessary to design a novel algorithm that can provide a high-quality solution within a short and predictable time. This paper proposes a lightweight attention-based scheduler called CoRaiS, and combines it with reinforcement learning to automatically learn a great policy that helps produce a high-quality scheduling decision in real time.
Figure 10: Matching-on-demand architecture of CoRaiS
VI-A Architecture of CoRaiS
CoRaiS adopts matching-on-demand (MoD) architecture that consists of two alignment modules (edge encoder and request encoder) and one matching module (context decoder), as presented in Fig. 10. The alignment modules are used to align specific features of heterogeneous edges through multi-dimensional information exchange. Specifically, the edge encoder embeds and aligns the performance information of edges through multi-head attention mechanism (MHA), and captures the service capacity of multi-edge system by max pooling; the request encoder has similar function with edge encoder, but works on the requests contexts, i.e. request encoder focuses on capturing and aligning the requirements of requests through MHA, and mastering the global request features by max pooling. The matching module (context decoder) associates the capacity of multi-edge computing system with the requirements of requests through aggregating edges embeddings and requests embeddings, and produces scheduling policies to realize edge matching based on the demands of request.
Edge encoder: It is a module to embed edge features. At the beginning of training, input features of edges $\mathbf{f}$ are initialized based on the current edge states. To be specific, the input features $\mathbf{f}_{q}$ of $e_{q}$ include (i) coordinates $(x_{q},y_{q})$; (ii) the coefficients of $\phi_{q}$ and replica numbers $\zeta_{q}$; (iii) workload evaluation vector $I_{q}$, and $I_{q}=(c_{q}^{le},c_{q}^{in},t_{q}^{in},b_{q}^{le},b_{q}^{in})$. The encoder computes initial $d_{h}$-dimensional edge embeddings $\mathbf{f}_{q}^{(0)}$ through a learnable linear projection with parameters $\mathbf{W}_{e}$ and $\mathbf{b}_{e}$: $\mathbf{f}_{q}^{(0)}=\mathbf{W}_{e}\mathbf{f}_{q}+\mathbf{b}_{e}$. Then the embeddings are updated through $L$ attention layers, which is motivated by Transformer [40] and Attention Model [41]. Each layer consists of two sublayers: a multi-head attention layer (MHA) and an edge-wise fully connected layer (FC). A skip-connection and batch normalization (BN) are also used at each sublayer. The operations are formulated as (16), where $\mathbf{f}_{q}^{(l)}$ denotes the produced edge embedding by layer $l\in\{1,..,L\}$.
$\begin{split}\mathbf{f}_{q}^{{}^{\prime}(l)}&=\mathbf{BN}^{l}(\mathbf{f}_{q}^{(l-1)}+\mathbf{MHA}_{e}^{l}(\{\mathbf{f}_{q}^{(l-1)}\}_{q=1}^{Q}))\\ \mathbf{f}_{q}^{(l)}&=\mathbf{BN}^{l}(\mathbf{f}_{q}^{{}^{\prime}(l)},\mathbf{FC}_{e}^{l}(\mathbf{f}_{q}^{{}^{\prime}(l)})\end{split}$ (16)
Request Encoder: It has similar architecture and operations with edge encoder to embed request features. But the learnable parameters are different. The initial feature $h_{z}$ of $r_{z}$ includes (i) coordinates of the source edge of $r_{z}$; (ii) input data size of $r_{z}$; (iii) coordinates of the predicted edge of $r_{z}$; (iv) output data size of $r_{z}$. The initial features $\mathbf{h}_{z}^{(0)}$ are initialized to a $d_{r}$-dimensional embeddings by linear projection as (17). Following that, $K$ attention layers are used to update request embeddings. The operations are presented as (18), where $MHA_{r}^{k}$ and $FC_{r}^{k}$ are the parameters that are used to embed requests features. The operations of $\mathbf{MHA}_{r}^{k}$ are similar with them in $\mathbf{MHA}_{e}^{l}$.
$\mathbf{h}_{z}^{(0)}=\mathbf{W}_{r}\mathbf{h}_{z}+\mathbf{b}_{r}$ (17)
$\begin{split}\mathbf{h}_{z}^{{}^{\prime}(k)}&=\mathbf{BN}^{k}(\mathbf{h}_{z}^{(k-1)}+\mathbf{MHA}_{r}^{k}(\{\mathbf{h}_{z}^{(k-1)}\}_{z=1}^{Z}))\\ \mathbf{h}_{z}^{(k)}&=\mathbf{BN}^{k}(\mathbf{h}_{z}^{{}^{\prime}(k)},\mathbf{FC}_{r}^{k}(\mathbf{h}_{z}^{{}^{\prime}(k)})\end{split}$ (18)
Context decoder: The system context comes from the edge embeddings $\{\mathbf{f}_{q}^{(L)}\}_{q=1}^{Q}$ and request embeddings $\{\mathbf{h}_{z}^{(K)}\}_{z=1}^{Z}$. Inspired by [42, 43, 44], this paper captures global features $\hat{\mathbf{f}}$ and $\hat{\mathbf{h}}$ by max pooling operation over $\{\mathbf{f}_{q}^{(L)}\}_{q=1}^{Q}$ and $\{\mathbf{h}_{z}^{(K)}\}_{z=1}^{Z}$, respectively. Then the new context embedding $c^{\prime}_{q}$ is produced by $\mathbf{MHA}_{c}$ which has $M$ heads. The computation process through $\mathbf{MHA}_{c}$ is shown as (19), where [.,.,.][.,.,.][ . , . , . ] is the horizontal concatenation operator, $d_{y}=\frac{d_{r}}{M}=\frac{d_{h}}{M}$, $\mathbf{c^{\prime}}_{q}$ is the embedding after single attention, and $\mathbf{c}_{q}$ is the final multi-head attention value for context embedding.
$\begin{split}&\mathbf{f}_{(c)q}=[\hat{\mathbf{f}},\hat{\mathbf{h}},\mathbf{f}_{q}^{(L)}]\\ &\mathbf{x}_{q}=\mathbf{W}_{c}^{x}\mathbf{f}_{(c)q};~{}~{}~{}\mathbf{y}_{z}=\mathbf{W}_{c}^{y}~{}\mathbf{h}_{z}^{(K)};~{}~{}~{}\mathbf{v}_{z}=\mathbf{W}_{c}^{v}~{}\mathbf{h}_{z}^{(K)}\\ &u_{qz}=\frac{\mathbf{x}_{q}^{T}\mathbf{y}_{z}}{\sqrt{d_{y}}};~{}~{}a_{qz}=\frac{e^{u_{qz}}}{\sum_{z=1}^{Z}e^{u_{qz}}}\\ &\mathbf{c^{\prime}}_{q}=\sum_{z=1}^{Z}a_{qz}\mathbf{v}_{z};~{}~{}\mathbf{c}_{q}=\sum_{i=1}^{M}~{}\mathbf{W}_{c,i}~{}\mathbf{c^{\prime}}_{q}^{i}\end{split}$ (19)
The $\mathbf{MHA}$ operation in context embedding is similar with it in edge encoder and request encoder, but replacing $\mathbf{f}_{(c)q}$ with $\mathbf{f}_{q}^{(l-1)}$ ($\mathbf{h}_{z}^{(k-1)}$) in edge encoder (request encoder).
Policy generation: The policy is generated by collecting importance of edges for one request $r_{z}$ as (20). $imp_{qz}$ denotes the importance of $e_{q}$ for $r_{z}$. $C$ is a constant. To evaluate the probability of edges getting the privilege, softmax is introduced over all edges for each request as (21), where $a_{qz}$ specifies the probability of $e_{q}$ responding to $r_{z}$.
$\begin{split}&\mathbf{px}_{q}=\mathbf{W}_{px}~{}\mathbf{c}_{q};~{}\mathbf{py}_{z}=\mathbf{W}_{py}~{}\mathbf{h}_{z}^{(H)};\\ &u_{qz}=\frac{\mathbf{px}_{q}^{T}\mathbf{py}_{z}}{\sqrt{d_{py}}};imp_{qz}=C\tanh(u_{qz})\end{split}$ (20)
$a_{qz}=\frac{e^{imp_{qz}}}{\sum_{q=1}^{Q}e^{imp_{qz}}}$ (21)
VI-B Training CoRaiS
The scheduling probability distribution $p_{\theta}(\pi)=\{a_{qz}\}_{q\in[Q],z\in[Z]}$ is produced by CoRaiS, from which we can sample a scheduling decision $\pi$. In order to train CoRaiS, we define the expectation of the maximum response time of all requests over edges as loss function: $\mathcal{L}(\theta|g)=\mathbb{E}_{p_{\theta}(\pi)}[L(\pi)]$. Given $\pi$, $L(\pi)=-\hat{u}^{\pi}$ is computed according to (22) and (23). Firstly, a local reward $u_{q}$ is estimated by (22), where $RL_{q}^{\pi}$ denotes the set of requests to be executed locally and $RT_{q}^{\pi}$ includes requests that are transferred from other edges to $e_{q}$, based on the decision $\pi$. $Y(\varpi_{m},e_{q})$ is a function to compute the transmission distance between the source edge $\varpi_{m}$ and the execution edge $e_{q}$ of $r_{m}$. $Y(\delta_{m},e_{q})$ obtains the transmission distance between the predicted edge $\delta_{m}$ and the execution edge $e_{q}$ of $r_{m}$. The connotation of each equation in (22) is similar with (7)-(13). Then the global reward is formulated as (23).
$\begin{split}\mu_{q}=&\frac{\sum_{r_{m}\in RL_{q}^{\pi}}\phi_{q}(f_{m})}{\zeta_{q}}+c_{q}^{le}\\ \eta_{q}=&\frac{\sum_{r_{m}\in RT_{q}^{\pi}}\phi_{q}(f_{m})}{\zeta_{q}}+c_{q}^{in}\\ \kappa_{q}=&\max\{\max\limits_{r_{m}\in RT_{q}^{\pi}}f_{m}~{}Y(\varpi_{m},e_{q}),~{}~{}t_{q}^{in}\}\\ \beta_{q}=&\max\{\max\limits_{r_{m}\in RT_{q}^{\pi}}u_{m}~{}Y(e_{q},\delta_{m}),b_{q}^{in}\}\\ u_{q}^{\pi}=&~{}-(\max(\mu_{q},\kappa_{q})+\max(\eta_{q}+\beta_{q},b_{q}^{le}))\end{split}$ (22)
$\hat{u}^{\pi}=\mathop{min}\limits_{e_{q}\in\mathcal{E}}~{}~{}u_{q}^{\pi}$ (23)
We use $\mathbf{S}$-samples batch reinforcement learning (RL) and gradient descent [45] to optimize $\mathcal{L}$, since the $\mathbf{S}$-samples batch gradient descent replacing the one-sample approximation in training realizes more accurate estimation of the policy, decreases training variance and speeds up convergence. Meanwhile, to encourage CoRaiS to sufficiently explore the huge search space, we add an extra entropy loss $H(\theta)$, computed as (24). Then the optimization function can be formulated as (25), where $D$ is the training data set, $C_{1}$ and $C_{2}$ are the coefficients.
$H_{\theta}(g)=-\sum_{z=1}^{Z}\sum_{q=1}^{Q}a_{qz}(\theta)\log a_{qz}(\theta)$ (24)
$\begin{split}A(\pi_{s})=&L(\pi_{s})-\frac{1}{S}\sum_{i=1}^{S}L(\pi_{i})\\ \mathcal{L}(\theta|\mathcal{D})=&\mathbb{E}_{g\sim D}(C_{1}\sum_{s=1}^{S}\log p_{\theta}(\pi_{s}|g)A(\pi_{s})\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}-C_{2}H_{\theta}(g))\\ \end{split}$ (25)
VI-C Decoding Strategies of CoRaiS
Two decoding strategies are proposed for CoRaiS to generate an effective scheduling decision.
  • Greedy decoding: according to the generated policy of CoRaiS, the best execution edge for each request is always selected. That is, for request $r_{z}$, its execute edge $e_{q}$ is selected by $q=\arg\max_{k}\{a_{kz}\}_{k=1}^{Q}$.
  • Sampling decoding: for request $r_{z}$, multiple edge selections can be sampled based on multinomial probability distribution over the policy $\{a_{qz}\}_{q=1}^{Q}$, and the best one is reported. A complete scheduling decision entails execution edge selections of requests in the multi-edge computing system.
VII Simulation Evaluation
To demonstrate that CoRaiS is able to learn a strong policy and give a real-time decision for the multi-edge scheduling problem, we designs three experiments: conventional test, generalization test, and characteristic validation. The conventional test refers to evaluating the performance of CoRaiS on a dataset that matches the same scale [^3] In this paper, the scale of dataset refers to the size of the multi-edge cooperative system, including the number of edges in the system and the number of requests submitted to it. as the one it was trained on. The generalization test refers to evaluating the performance of CoRaiS on a dataset that has large scales than the one it was originally trained on. The characteristic validation is used to assess whether CoRaiS can perceive the workload and heterogeneity of edges, and autonomously implement load balancing.
TABLE III: Normal Test Results
CoRaiS is trained and tested on the same small-scale instances $\{(10,5,50),(10,5,100),(20,10,50),(20,10,100)\}$.
Gap-M, Time-M and Time-S: the lower the better.
(10, 5, 50) (10, 5, 100) (20, 10, 50) (20, 10, 100)
Methods Gap-M Time-M Time-S Gap-M Time-M Time-S Gap-M Time-M Time-S Gap-M Time-M Time-S
Gurobi(10s) 1.0000 0.4294 1.1454 1.0000 1.3126 1.9815 1.0000 0.1970 0.0170 1.0000 0.5791 1.3249
Local 1.9602 0.0049 0.0008 2.7110 0.0047 0.0006 1.5214 0.0085 0.0008 2.0253 0.0081 0.0015
Predicted 1.9752 0.0090 0.0015 2.7182 0.0123 0.0013 1.5268 0.0125 0.0011 2.0256 0.0155 0.0028
Random(1) 2.0540 0.0049 0.0008 2.7707 0.0048 0.0005 1.5706 0.0086 0.0008 2.0943 0.0082 0.0015
 Random(100) 1.4199 0.0053 0.0008 1.9525 0.0053 0.0006 1.1211 0.0091 0.0008 1.4544 0.0089 0.0016
Random(1k) 1.2962 0.0066 0.0011 1.7691 0.0079 0.0011 1.0624 0.0104 0.0008 1.3342 0.0112 0.0020
FC1-CoRaiS(greedy) 1.4953 0.0135 0.0019 1.5857 0.0132 0.0016 1.4443 0.0171 0.0015 1.5014 0.0163 0.0029
FC2-CoRaiS(greedy) 1.1738 0.0151 0.0034 1.3871 0.0145 0.0015 1.0504 0.0175 0.0025 1.1479 0.0175 0.0031
FC3-CoRaiS(greedy) 1.1802 0.0118 0.0017 1.3422 0.0115 0.0012 1.0345 0.0150 0.0018 1.1279 0.0147 0.0025
CoRaiS(greedy) 1.0476 0.0159 0.0022 1.0823 0.0155 0.0015 1.0182 0.0193 0.0016 1.0409 0.0187 0.0031
FC1-CoRaiS(100) 1.2465 0.0138 0.0019 1.3366 0.0134 0.0016 1.2552 0.0170 0.0015 1.2291 0.0165 0.0029
FC2-CoRaiS(100) 1.0544 0.0150 0.0034 1.2422 0.0146 0.0015 1.0072 0.0181 0.0025 1.0266 0.0179 0.0031
FC3-CoRaiS(100) 1.0566 0.0120 0.0017 1.1949 0.0117 0.0012 1.0038 0.0152 0.0018 1.0259 0.0150 0.0025
CoRaiS(100) 1.0160 0.0157 0.0023 1.0353 0.0154 0.0016 1.0000 0.0192 0.0016 1.0051 0.0185 0.0031
FC1-CoRaiS(1k) 1.1957 0.0138 0.0019 1.2757 0.0135 0.0016 1.2229 0.0166 0.0015 1.1805 0.0166 0.0029
FC2-CoRaiS(1k) 1.0426 0.0151 0.0034 1.2167 0.0146 0.0015 1.0030 0.0183 0.0025 1.0193 0.0179 0.0031
FC3-CoRaiS(1k) 1.0446 0.0122 0.0017 1.1750 0.0117 0.0012 1.0019 0.0156 0.0018 1.0178 0.0151 0.0025
CoRaiS(1k) 1.0113 0.0159 0.0022 1.0274 0.0153 0.0018 1.0000 0.0192 0.0018 1.0036 0.0185 0.0032
Gurobi(0.02s) (0.8%) 0.0772 0.0033 (0%) - - (0%) - - (0%) - -
Gurobi(0.05s) (78.61%) 0.1064 0.0093 (3.2%) 0.1760 0.0162 (2.81%) 0.1713 0.0092 (0%) - -
Gurobi(1s) 1.0001 0.2489 0.3041 1.0015 0.5938 0.4399 1.0000 0.2115 0.0176 1.0005 0.4231 0.1432
Gurobi(5s) 1.0000 0.3963 0.8883 1.0000 1.1848 1.4795 1.0000 0.1991 0.0172 1.0000 0.4875 0.6733
VII-A Experiment Descriptions
Instance generation   The training and testing datasets are generated as the same rules. Given the number of edges in the whole multi-edge cooperative computing system, denoted as $N$, the number of edges in an individual control region $Q(Q<N)$ and the number of requests $Z$, for each edge $e_{q}(q\in\{1,...,N\})$, the coordinates are randomly sampled under the uniform distribution in $(0,1)^{2}$; the supported maximum service replica number $rn_{q}$ is randomly sampled in $\{1,2,3,4\}$. $C_{t}=3$. The functional relationship between computing time and the size of input data packets is modeled as linear functions [^4] Without loss of generality, we model the relationship using the linear functions. But in practical, other relationships, such as quadratic function, are allowed to train CoRaiS. for all edges during simulations. Moreover, to represent the heterogeneity across edges, different coefficients are randomly sampled from a uniform distribution within the range $(0,1)$.
To make CoRaiS can learn to adopt to any initial system-level states, we randomly generate some requests as backlogs for each edge while generating training datasets and testing datasets. The numbers of backlogs in $Q_{q}^{le}$ and $Q_{q}^{in}$ of edge $e_{q}$ are randomly sampled from $(0,100)$. For any backlog $r_{x}$ in $Q_{q}^{le}$, its input data size $f_{x}$ is uniformly sampled from $(0,1)$, output data size $u_{x}$ is uniformly sampled from $(0,0.1)$, and its predicted edge $\delta_{x}$ is sampled from $[N]$. For any backlog $r_{k}$ in $Q_{q}^{in}$, its source edge $\varpi_{k}$ is sampled from $[Q]-\{q\}$, its predicted edge $\delta_{k}$ is sampled from $[N]$, and its input and output data sizes are uniformly sampled from $(0,1)$ and $(0,0.1)$. With these backlogs, the service-oriented workload evaluation is carried out as (1)-(3).
For any new request $r_{z}$ that will be scheduled in current period, its source edge $\varpi_{z}$ is sampled from $[Q]$, its predicted edge $\delta_{z}$ is sampled from $[N]$, and the related input data size $f_{z}$ and output data size $u_{z}$ are uniformly sampled from $(0,1)$ and $(0,0.1)$, respectively.
Hyperparameters   Learnable parameters are initialized as Uniform($-1/\sqrt{d},1/\sqrt{d}$), with $d$ is the input dimension; learning rate $lr=1e-5$; batch size is 128, $S=64$ while using $S$-samples batch RL; $C_{1}=10$ and $C_{2}=0.5$. The edge encoder and request encoder have $L=5$ and $K=3$ attention layers, respectively. $\mathbf{MHA}$ and $\mathbf{FC}$ in two embedding modules have same structure. $\mathbf{MHA}$ has 8 heads, $\mathbf{FC}$ has one hidden sublayer with dimension 512 and ReLu activation.
CoRaiS is trained from 40000 batches on the device with 2×2$\times$2 ×Intel(R) Xeon(R) Gold 5215 CPU that can provide 40 processors and 2×2$\times$2 ×NVIDIA GTX 2080Ti, and CoRaiS uses Adam optimizer, PyTorch 1.11 framework and python 3.8 on Ubuntu 18.04.
Baselines   It must be mentioned that as our best knowledge, CoRaiS is the first artificial intelligence based work that breaks many assumptions of prior works, deals with a multi-edge computing system where every edge is allowed to receive requests from clients, and optimizes an objective function that minimize the responding time over all requests. In order to evaluate the performance of CoRaiS, the results generated by CoRaiS are compared with that produced by the exact solver Gurobi. Gurobi is able to calculate exact solutions. Meanwhile, we propose three heuristic approaches (Local, Random and Predicted) as the lightweight scheduling baselines in the scenarios. Three learning-based baselines are designed as well.
  • Exact baseline: Gurobi is one of the state-of-the-art solver for integer linear programming problems. However, Gurobi may fall into long-term calculation due to the huge search space of multi-edge scheduling. Therefore, it is necessary to give a computation time limitation $(x~{}s)$.
  • Lightweight heuristic baselines: two algorithms are used to provide heuristic baselines. (i) Local: executing all requests at their source locations. (ii) Random: randomly sampling the execution edges for all requests. It is allowed to sample multiple times and report the best one. (iii) Predicted: executing all requests at their predicted edges, but if the predicted edges of some requests are out of the control region, the requests will be executed at their source edges.
  • Learning-based baselines (ablation studies for aligning modules of CoRais): three synthetic neural network models are used to illustrate the effectiveness of CoRaiS. (i) FC1-CoRaiS: replacing the multi-head attention aligning mechanism of edge encoder in CoRaiS with the multi-layer perceptron, and maintaining the same number of neuron parameters. (ii) FC2-CoRaiS: adopting the similar MoD architecture with CoRaiS, but using the multi-layer perceptron to replace the multi-head attention aligning mechanism in request encoder. (iii) FC3-CoRaiS: adopting MoD structure, but without aligning mechanisms in both edge encoder and request encoder, only using multi-layer perceptron to embed edges’ features and requests’ features. Since FC1-CoRaiS, FC2-CoRaiS and FC3-CoRaiS adopt the similar MoD architecture and have the same input/output with CoRaiS, they are able to use the same decoding strategies with CoRaiS as well.
TABLE IV: Generalization Test Results on Medium-scale Instances.
CoRaiS is trained on small-scale instances (10, 5, 100) and directly applied into medium-scale instances. $(x\%)$ refers to $x\%$ of the instances are solved.
Gap-M, Time-M and Time-S: the lower the better.
(20, 10, 200) (60, 30, 400) (100, 50, 600) (100, 50, 800)
Methods Gap-M Time-M Time-S Gap-M Time-M Time-S Gap-M Time-M Time-S Gap-M Time-M Time-S
Gurobi(10s) 1.0000 2.0173 3.0761 1.0000 6.1590 1.1698 (97.60%) 16.7801 1.2889 (58.00%) 20.0647 0.7842
Local 3.0285 0.0088 0.0008 2.4070 0.0243 0.0019 2.2841 0.0394 0.0028 2.4602 0.0394 0.0024
Predicted 3.0860 0.0265 0.0018 2.4735 0.0620 0.0032 2.3190 0.0968 0.0047 2.4820 0.1145 0.0042
Random(1k) 1.9392 0.0147 0.0013 1.6445 0.0362 0.0024 1.5753 0.0574 0.0034 1.7049 0.0608 0.0034
CoRaiS(greedy) 1.0739 0.0206 0.0014 1.0139 0.0373 0.0020 0.9822 0.0533 0.0030 0.9400 0.0523 0.0025
CoRaiS(100) 1.0246 0.0201 0.0016 1.0002 0.0362 0.0021 0.9672 0.0523 0.0033 0.8999 0.0520 0.0028
CoRaiS(1k) 1.0202 0.0202 0.0015 1.0002 0.0357 0.0022 0.9672 0.0519 0.0032 0.8999 0.0514 0.0028
Gurobi(0.05s) (0%) - - (0%) - - (0%) - - (0%) - -
Gurobi(1s) 1.0062 0.9511 0.2605 (0%) - - (0%) - - (0%) - -
Gurobi(5s) 1.0006 1.4622 1.5466 1.0032 6.0602 0.8031 (1.2%) 12.4700 0.4302 (0%) - -
Performance indexes   Three indexes are used to evaluate performance.
  • Gap-M: average quality difference of solutions that are generated by CoRaiS and other baselines, compared with the solution generated by Gurobi(10s) in the same instance. Solutions from Gurobi(10s) are considered as the optimal benchmark in the simulations. The gap is computed by (26). $\aleph$ includes CoRaiS and other baselines. $\pi$ signifies the best solution obtained from the specific approach $b$ ($b\in\aleph$). $\hat{\pi}$ refers to the best solution generated by $Gurobi(10s)$. $L(\pi)$ refers to the predicted response time for all requests in the multi-edge cooperative computing system, and $L(\pi)$ is computed by (23). Gap-M is critical because the multi-edge cooperative computing system also require the real-time approach to produce a high-quality solution.
    $gap_{b}=\frac{L(\pi|b)}{L(\hat{\pi}|Gurobi(10s))},~{}~{}\forall b\in\aleph$ (26)
  • Time-M: average time taken to make scheduling decisions. Time-M is an important index to estimate whether the methods can provide effective decision in time.
  • Time-S: the standard variance of computing times to make scheduling decisions. Time-S can evaluate the stability of decision-making time.
TABLE V: Generalization Test Results on Lager-scale Instances
CoRaiS is trained on small-scale instances (10, 5, 100). Cost-M, Gap-M and Time-M: the lower the better.
(200, 100, 2k) (600, 300, 4k) (1000, 500, 6k) (2k, 1k, 10k) (2k, 1k, 15k)
Methods Cost-M Gap-M Time-M Cost-M Gap-M Time-M Cost-M Gap-M Time-M Cost-M Gap-M Time-M Cost-M Gap-M Time-M
Gurobi(5s) (0%) - - (0%) - - (0%) - - (0%) - - (0%) - -
Gurobi(10s) (3.1%) - - (0%) - - (0%) - - (0%) - - (0%) - -
CoRaiS(1k) 5.1795 1.0000 0.0600 5.8071 1.0000 0.3783 6.0991 1.0000 0.9563 7.1192 1.0000 3.1611 8.0548 1.0000 4.6019
Local 16.6937 3.2390 0.0403 14.5508 2.5340 0.1285 14.1081 2.3484 0.2129 13.7089 2.0241 0.6159 16.7721 2.2784 0.4401
Predicted 17.0419 3.3059 0.1293 14.5757 2.5398 0.3452 14.1772 2.3591 0.4828 13.7813 2.0354 1.2682 16.7845 2.2807 1.1119
Random(100) 13.0677 2.5316 0.0441 12.0275 2.0924 0.1426 11.8956 1.9807 0.2227 11.6695 1.7260 0.6598 14.4036 1.9613 0.6160
Random(500) 12.4988 2.4216 0.0573 11.6141 2.0199 0.2261 11.5509 1.9227 0.5283 11.3679 1.6806 1.7379 14.0588 1.9127 2.4922
Random(1k) 12.3202 2.3865 0.0796 11.4478 1.9912 0.4145 11.4235 1.9008 1.0049 11.2717 1.6671 3.2203 13.9004 1.8908 4.5946
CoRaiS(10) 5.2763 1.0186 0.0485 6.0588 1.0396 0.1547 6.4901 1.0604 0.2459 7.9634 1.1092 0.7364 9.2483 1.1436 0.6991
CoRaiS(100) 5.2053 1.0049 0.0597 5.9007 1.0139 0.1562 6.1954 1.0147 0.2724 7.4796 1.0441 0.7771 8.5213 1.0547 0.8522
CoRaiS(500) 5.1847 1.0009 0.0510 5.8259 1.0027 0.2175 6.1188 1.0035 0.5298 7.1911 1.0093 1.7753 8.1929 1.0183 2.6391
TABLE VI: Characteristics Validation Results
CoRaiS is trained on scheduling problems (10, 5, 100).    CoRaiS(1k) is used to sample near-optimal solutions. LB: Load Balancing; WP: Workload Perception; HA: Heterogeneity Awareness; EReqN: the number of requests that are executed at each edges; LCost: the corresponding response time.
LB WP HA
No. EReqN LCost EReqN LCost EReqN LCost
A 21.5950 4.1647 19.7030 14.7180 11.9850 4.2231
B 19.6170 4.1958 20.2250 14.9676 14.8660 4.2165
C 19.5190 4.1825 20.0240 14.8704 21.5160 4.0770
D 19.6090 4.1935 20.0240 14.8709 25.8380 3.6357
E 19.6600 4.1953 20.0240 14.8709 25.7950 2.4347
VII-B Results Analysis
VII-B1 Analysis of Conventional Test Results
CoRaiS is a lightweight model that has about 4 million learnable parameters. We trained CoRaiS on four scales, i.e., $(N,Q,Z)\in\{(10,5,50),(10,5,100),(20,10,50),(20,10,100)\}$ where $(N,Q,Z)$ denotes the number of edges in the whole multi-edge cooperative computing system, the number of edges in an individual control region $Q$ and the number of requests $Z$. We test the performance of learned policies on the same scale problems, and the results are shown in Table III.
(i) Gurobi can quickly obtain the optimal solutions for small-scale problems, and the average time costs for $(N,Q,Z)\in\{(10,5,50),(20,10,50),(20,10,100)\}$ are less than $1s$. However, as the problem scale becomes larger, such as when $(N,Q,Z)=(10,5,100)$, the solving time increases significantly, even sometimes, Gurobi spends $10s$ but only get a sub-optimal solution.
(ii) Lightweight heuristic approaches (Local, Predicted and Random) take near-zero time to give a solution, but the quality is usually far away from the optimum.
(iii) The solving time taken by CoRaiS(greedy, 100, 1k) is close to $0.02s$ with low variance $0.003s$, which is less than Gurobi in both Time-M and Time-S. The gap closing to 1 shows CoRaiS successfully learns an efficient policy that produces high-quality solutions. Therefore, CoRaiS has potential to provide real-time and high-quality decision to support efficient operations of the multi-edge cooperative computing system.
(iv) CoRaiS and other three learning-based models (FC1-CoRaiS, FC2-CoRaiS and FC3-CoRaiS) adopt the same decoding strategies (greedy, sampling). The comparison results show that the two alignment mechanisms of edge/request features play important roles in promoting policy learning.
(v) Because CoRaiS spends near $0.02s$ on solving the problems, we explore the performance of Gurobi(0.02s) as well. The experimental results illustrate that it is very difficult for Gurobi to solve the problem within such a short time, even to calculate an approximate solution. Then we relax the time constraints to $0.05s$, $1s$ and $5s$. The results generated by Gurobi(0.05 s) show that double relaxation does not improve performance much. The proportion of problems that are not solved approximately is still high. Only some simple problems can obtain near-optimal solutions. The results generated by Gurobi(1s) and Gurobi(5s) show 50X and 200X relaxation help a lot, since all problems get their sub-optimal solutions. However, the average decision-making time is too long for a real-time computing system, and the stability of time usage also performs not well.
VII-B2 Analysis of Generalization Test Results
We train CoRaiS on instances under problem scale setting $(10,5,100)$. The learned model is directly applied into medium-scale and large-scale instances. The results are presented in Table IV and Table V, respectively.
In the medium-scale instances, the number of edges are limited to be more than 10 but less than 100, while the number of requests are constrained to several hundred levels. We present the results on four medium scales in Table IV, including $(20,10,200),(60,30,400),(100,50,600)$ and $(100,50,800)$. Gurobi(10s) takes a long time to compute a good solution, and its time cost increases as the problem scale becomes larger, i.e. from $2s$ for $(20,10,200)$ to $20s$ for $(100,50,800)$. However, as the scales become too large, Gurobi(10s) cannot solve the problem within the limited computing time, such as instances on $(100,50,600)$ and $(100,50,800)$. Compared with Gurobi(10s), CoRaiS costs a shorter time $(<0.05s)$ to obtain a high-quality near-optimal solution by sampling, and its time cost does not increase significantly, from $0.02s$ for $(20,10,200)$ to $0.05s$ for $(100,50,800)$, even though the problem scale becomes 20X larger. We force Gurobi to provide solutions within $0.05s$ to check whether it can get similar performance with CoRaiS. The results presented in Table IV show that Gurobi can hardly solve large-scale problems within a short time. Then we relax the time limitation to 20X and 100X, i.e. $1s$ and $5s$, the comparison results show that only a few large-scale problems can be solved. Lightweight heuristic approaches are also used as baselines. The comparison results show that even though CoRaiS has a decision-making time twice that of heuristic approaches, it can produce much better assignment decisions to realize shorter response time over all requests.
Furthermore, we conduct generalization testing on large-scale instances. The number of edges in large-scale instances is set to up to 1k, while the number of requests ranges from several thousand to 15k. Table V displays the results on five scales, including (200, 100, 2k), (600, 300, 4k), (1k, 500, 6k), (2k, 1k, 10k) and (2k, 1k, 15k). To facilitate the evaluation of different algorithms’ performance, we introduce a new metric, Cost-M, estimating the average response time over requests, measured in seconds. In the table, we observe that Gurobi(5s) and Gurobi(10s) struggle to handle the large-scale instances. Therefore, we select CoRaiS(1k) as the baseline and compare its results with lightweight heuristics. It’s evident that decision-making time increases with problem scales for all scheduling approaches. Lightweight heuristic approaches, especially Local, enable faster decision-making. However, their decision quality is significantly inferior to that of CoRaiS(1k). The Cost-M of lightweight heuristics is usually more than twice that of CoRaiS(1k). Additionally, to address the need for swift and effective decision-making in specific cases, we relax the sampling times of CoRaiS to 10, 100, and 500. The results show that with fewer samples, CoRaiS can speed up decision-making while ensuring more effective scheduling than baseline methods.
VII-B3 Analysis of Characteristic Validation Results
(i) Load balancing (LB): We design five homogeneous edges and push same backlogs on them, so that the initial system-level state of edges are same, and the time of edges responding to backlogs satisfies $b_{E}=b_{D}=b_{C}=b_{B}=b_{A}$. Then we conduct 10k experiments on the system. In each experiment, 100 same requests are submitted to $e_{A}$, and CoRaiS is used to provide a scheduling solution. The results are presented in Table VI(LB). The number of requests executed at five edges are approximately equal, and the response time are very close. Therefore, we claim that CoRaiS learned to optimize scheduling through load balancing without manual intervention.
(ii) Workload perception (WP): We design five homogeneous edges and induce differences in response time to backlogs by pushing different numbers of requests to each edge, so that the initial system-level state of edges are different. The time of edges responding to backlogs satisfies $b_{E}\leq b_{D}\leq b_{C}\leq b_{B}<b_{A}$. Then 10k experiments submitting 100 requests to $e_{A}$ are conducted. CoRaiS makes decisions to schedule requests. As shown in Table VI(WP), the number of requests to be dispatched to each edge is different and follows the order $n_{E}\geq n_{D}\geq n_{C}\geq n_{B}>n_{A}$ on average. Meanwhile, there is no significant difference in response time after scheduling. Therefore, we claim CoRaiS is able to perceive workload of edges while scheduling.
(iii) Heterogeneity awareness (HA): We design five heterogeneous edges, that is, the computation time estimation functions of edges are different. We arrange the computing performance of edges in the order of $E>D>C>B>A$. Then we equalize the response times of the edges to backlogs by adjusting the number of requests. After that, 10k experiments are conducted. In each experiment, 100 same requests are submitted to $e_{A}$, and CoRaiS is used to provide a scheduling solution. According to the results in Table VI(HA), the more powerful edge serves more requests, and the corresponding response time of all edges are very close. Therefore, we claim CoRaiS is able to recognize heterogeneity while cooperating multi-edges.
VIII Prototype Evaluation
Figure 11: The linear relationship between running time and the size of data packets for Model Soups [37] on the five heterogeneous devices (Edge A, Edge B, Edge C, Edge D, Edge E). Points are actual statistical data, while lines are function relationships fitted based on the data.
VIII-A Environment Description
Prototype experiments are conducted to illustrate the performance of proposed CoRaiS. Through out the experiments, five heterogeneous servers are deployed as edge computing nodes, and the servers are networked using a dedicated VPN. The configurations and the computation time estimation functions (Model Soups[37] oriented) of the servers are shown in Fig. 11. The number of service replicas deployed on the edges are (2, 4, 3, 3, 1). Three levels of requests are submitted to the edges, comprising small-size requests with 2.6MB of data, medium-size requests with 7.9MB, and large-size requests with 18.6MB. Five-scale experiments are established, i.e., the number of requests involves 50, 100, 200, 500, and 1000. In each experiment, the distribution percentages for these request sizes are 50%, 30%, and 20%, respectively. The primary index collected is the physical response time, considered the most crucial metric.
Two kinds of experiments are carried out, including equal-distribution testing and shifted-distribution testing.
  • Equal-distribution testing: Each request’s source edge and predicted edge are uniformly sampled from the five available edges, while its execution edge is determined through scheduling algorithms. The distribution setting for requests aligns with the data generation method used during the training of CoRaiS.
  • Shifted-distribution testing: The probabilities of each request being submitted to the five edges are predefined, i.e. [50%, 20%, 20%, 5%, 5%], which makes the initial distribution of requests on edges be quite different from the uniform distribution. The probabilities of each request’s predicted edge is predefined as well, that is, [10%, 10%, 40%, 25%, 15%]. The execution edge of each request is decided by scheduling algorithms.
To enable real-time scheduling of requests, lightweight scheduling algorithms were utilized in the experiments. These include CoRaiS [^5] CoRaiS is trained on scheduling problems (10, 5, 100). with a sampling decoding strategy, denoted as CoRaiS(1k), as well as three heuristic approaches capable of making decisions within 0.1 seconds, i.e. Local, Predicted, and Random(1k).
VIII-B Results Analysis
For equal-distribution testing, the practical response times of requests are summarized in Table VII. From the table, it is evident that CoRaiS(1k) consistently facilitates more effective scheduling decisions, resulting in faster response times for requests.
For shifted-distribution testing, the experimental results are summarized in Table VIII. The findings outlined in Table VIII reveal that CoRaiS is adaptable to request distributions different from those it was trained on, enabling more effective decisions to accelerate response times for requests.
TABLE VII: Practical response time of requests while establishing equal-distribution testing (measured in seconds)
ReqNum CoRaiS(1k) Local Predicted Random(1k)
50 6.361 11.206 8.223 7.985
100 13.503 14.375 15.891 15.101
200 26.613 33.124 31.087 35.854
500 67.972 72.965 77.099 80.369
1000 130.385 141.639 146.179 157.577
TABLE VIII: Practical response time of requests while establishing shifted-distribution testing (measured in seconds)
ReqNum CoRaiS(1k) Local Predicted Random(1k)
50 7.454 12.834 8.624 8.187
100 13.743 22.908 17.056 16.888
200 27.604 46.715 30.434 31.749
500 67.630 115.215 94.336 78.333
1000 131.461 224.968 176.093 157.254
IX Conclusion
To facilitate the effective cooperation among edges, this paper firstly introduces the system-level state evaluation model to shield edges heterogeneity in hardware configuration and redefine the service capacity at edges. Following that, the territorial schedulers are introduced into the service-oriented multi-edge cooperative computing system, and the multi-edge scheduling problem is formulated to an integer linear programming formulation to inspire the design of multi-edge cooperative algorithms. After that, CoRaiS, a learning-based lightweight real-time scheduler, is proposed to minimize the response time over all distributed arriving requests. The experimental results demonstrate that CoRaiS successfully learned a strong policy to make high-quality scheduling in real time, irrespective of request arrival patterns and system scales.
The proposed cooperation models and algorithms still have limitations in supporting multi-edge cooperation at a global level due to the lack of support for multi-scheduler cooperation. In the future, we will continue studying multi-edge cooperative computing and plan to design a hierarchical interoperability model based on current research progresses. The model will categorize schedulers into medium-level schedulers and a top-level scheduler. Medium-level schedulers can cooperate with each other under the control of their higher-level scheduler. With the improved interoperability model, scheduling problems will be reformulated from the different perspectives of medium-level schedulers and the top scheduler. Scheduling algorithms will also be designed with consideration of the accumulative decision-making time delay across the hierarchy of schedulers. The updated cooperation models and algorithms aim to ensure the effective operation of a scalable multi-edge cooperative computing system, regardless of the scheduling levels and the number of edges and requests in the system.
Acknowledgement
We are grateful to New Lossless Data Center at Purple Mountain Laboratories for generously providing sufficient computing power for model training and inferences.